OpenStudy (peachytea):

-2 cos (-300 degrees)

OpenStudy (hyuna301):

we can also write 300 degree is radians = 300 . pi/180 = 5.pi/3 now, cos ( 5pi/3) can also be written as.... cos( 2pi - pi/3) which is also equals to.... cos(pi/3) and cos(pi/3) = 1/2 also, cos(-theta) = cos(theta) now, the given question , -2 cos(-300 deg.) = -2 cos(300 deg.) = -2. 1/2 = -1